OpenStudy (anonymous):

Jason's savings account has a balance of $2179. After 5 years, what will the amount of interest be at 6% compounded quarterly?

OpenStudy (anonymous):

u must be working it out

OpenStudy (anonymous):

\[A=P \left( 1+\frac{ R }{ 100 } \right)^{T}\] \[A=2179\left( 1+\frac{ 6 }{4*100 } \right)^{5*4}=2179\left( \frac{ 203 }{ 200 } \right)^{20}\] Solve it.

OpenStudy (anonymous):

compound interest=A-P

OpenStudy (anonymous):

$746.80

OpenStudy (anonymous):

this what i got

OpenStudy (anonymous):

i calculated, it is 755.80 $

OpenStudy (anonymous):

ok help me with this one

OpenStudy (anonymous):

Nora's savings account has a balance of $3979. After 4 years, what will the amount of interest be at 12% compounded semiannually?

OpenStudy (anonymous):

i got this $2367.92

OpenStudy (anonymous):

\[A=3979\left( 1+\frac{ 12 }{ 2*100 } \right)^{4*2}=3979*\left( \frac{ 53 }{ 50 }\right)^{8}=6341.92\] Interest=6341.92-3979=2362.92 $

OpenStudy (anonymous):

whats your email?

OpenStudy (anonymous):

sshayer52@gmail.com

OpenStudy (anonymous):

ok ill email u later tonight for help